# Entanglement teleportation along a regenerating hamster-wheel graph state

### v1

@@ Abstract
We scheme an efficient and reusable approach to quantum teleportation that allows cyclic teleportation of a two-qubit graph state around a quantum hamster wheel—a ring of qubits entangled as a one-dimensional line prepared on the 20-qubit Quantinuum H1-1 ion-trap quantum processor. The qubits on the ring are periodically measured and reused to achieve a teleportation depth that exceeds the total number of available qubits in the quantum processor. Using the outcomes measured during teleportation, we calculate and apply byproduct operators through dynamic circuits to correct local transformations induced on the teleported state. We evaluate the quality of teleportation by tracing the preserved entanglement and fidelity of the teleported two-qubit graph state from its density matrix. In the real-machine experiments, we demonstrate that 58% of the entanglement of the teleported state is sustained with a measured two-qubit negativity of $0.291\pm 0.018$ after three complete revolutions around the hamster wheel, or equivalently, after hopping across 56 qubits. By performing teleportation along a regenerating graph state, our work is a step forward in demonstrating the feasibility of measurement-based quantum computation.
@@
In the following contents, we cover the experimental setups and methodologies for implementing quantum hamster wheels, along with the mathematics of teleportation. Next, we introduce the concepts of negativity and fidelity as the two useful indicators of the amount of entanglement and the success of teleportation. Finally, we report and summarise the results obtained from the execution of quantum hamster wheels both on the emulator and on the real machine, which sustained the entanglement of the two-qubit state of high quality after 75 hops and 56 hops in teleportation, respectively.
@@
Figure 3: Polar plots of the teleported two-qubit graph state showing (a) negativity and (b) fidelity on the radial axis versus the number of teleportation hops in the angular axis sampled from the H1-1 emulator averaged over four trials. The three data points marked in red represent results obtained from the real H1-1 quantum computer for 9, 18 and 56 hops with respective negativities $0.459\pm 0.009,0.388\pm 0.014,0.291\pm 0.018$ and fidelities $0.958\pm 0.009,0.887\pm 0.014,0.791\pm 0.018$, where experiments were conducted on 21 March, 25 March and 24 April 2024 respectively. Assuming the distribution is Gaussian, the half-width of the shaded strip is two times the standard error of the sampled negativity or fidelity to represent 95% confidence interval. On the other hand, the data obtained from real H1-1 quantum computers are only sampled in a single trial due to limited resources. Therefore, we use the technique of bootstrapping to resample the data from the measured sample distribution for 200 trials. The error bars represent 95% confidence interval of the bootstrapped negativities or fidelities. All results are mitigated with REM as indicated in Teleportation quality witnesses. Due to the feedforward gates applied based on mid-circuit measurements in the dynamic circuit approach, REM is applied to only the two qubits of the final teleported state.
@@
In Figure 3a and Figure 3b, we present the negativities and fidelities recorded from the real machine (red-marked data points) and the emulator (blue data), plotted in polar plots against the number of hops that the two-qubit graph states were teleported across using the teleportation strategies. The emulated data is averaged over four trials. Due to limited access to computational resources, experiments on the H1-1 quantum computer were conducted for only three scenarios of 9, 18, and 56 hops in a single trial of the QST experiment with 1000 shots for each QST circuit. We bootstrap 200 samples from each QST basis measurement for both negativities and fidelities. The error bars of the data obtained from the H1-1 quantum computer of size $\epsilon$ take the half-range of the 95% confidence interval of the bootstrapped negativities or fidelities, such that $\epsilon=\frac{1}{2}{\left(\left\lfloor a_{0.975N}\right\rfloor-\left\lceil a_{0.025N}\right\rceil\right)}$, where $\{a_{n}\},n=1,2,\cdots,N$ is the sequence of bootstrapped negativities or fidelities ordered from lowest to highest and $N=200$ is the sample size. For both the negativity and fidelity, we have used REM, the nearest physical probability vector, and the nearest physical density matrix algorithm as previously outlined to mitigate classical readout errors. After all, from the H1-1 quantum computer we obtain the negativity of $0.459\pm 0.009,0.388\pm 0.014,0.291\pm 0.018$ and fidelity of $0.958\pm 0.009,0.887\pm 0.014,0.791\pm 0.018$ corresponds to 9, 18 and 56 hops respectively. For the data obtained from the emulator, we report a clear trend that the entanglement of the teleported state decreases gradually, almost linearly, as the number of hops increases. This trend is closely in agreement with the data points obtained from the real device without significant deviations, as well as the trend of the two-qubit gate fidelity plot in the H1 model product data sheet [16]. Therefore, it reinforces the reliability of the H1-1 emulator to approximate the corresponding device noise efficiently. From the plot, we also find that both dynamic circuit and post-selection approaches exhibit very similar behaviour regarding the amount of entanglement. This observation agrees with our expectations that the two approaches should only introduce little differences in the negativity and fidelity, given that their circuit depth differs by at most three gates. On the other hand, we have shown a distinct utility of the post-selection approach against dynamic circuits presented in previous work [6] using IBM quantum computers. This was a direct consequence of the accumulated latency for processing each feed-forward mid-circuit measurement sequentially and the relatively high level of noise in IBM quantum computers. From these results, we confirm that even after 56 hops, the negativity of the state teleported using a dynamic circuit can be preserved up to $0.291\pm 0.018$. When tested using the emulator, it still maintains a value of $0.251\pm 0.001$ after 75 hops which is equivalent to 4 full cycles. Therefore, it is reasonable to expect that the actual limit of the number of hops the two-qubit state can teleport across using H1-1 quantum computer while preserving entanglement is beyond 100.
@@
Discussion.—In summary, we demonstrated the teleportation of two-qubit graph states cyclically on a ‘quantum hamster wheel’. By making use of the reset gates, we were able to reuse the measured qubits and perform teleportation hops more than the total number of qubits available in the Quantinuum 20-qubit H1-1 quantum computer with ion-trap architecture. Despite limited real machine resources, in the emulation, we still showed that the entanglement of the two-qubit state was preserved with the negativity of $0.291\pm 0.018$ after 56 hops in teleportation on the real quantum processor and $0.251\pm 0.001$ after 75 hops on the corresponding emulator with the device-noise model. We expect this figure to be extended above 100 without the break-off of entanglement. Such performance benchmarks the utility of ion-trap quantum processors in achieving high-fidelity two-qubit gates and long coherence times that are more robust against perturbations. Moreover, since the teleportation protocol with dynamic circuits shares a similar principle and process as the single-qubit measurement-based quantum computation, these results also serve as a feasibility demonstration of measurement-based quantum computation.

### v2

@@ Abstract
We scheme an efficient and reusable approach to quantum teleportation that allows cyclic teleportation of a two-qubit graph state around a quantum hamster wheel—a ring of qubits entangled as a one-dimensional line prepared on the 20-qubit Quantinuum H1-1 ion-trap quantum processor. The qubits on the ring are periodically measured and reused to achieve a teleportation depth that exceeds the total number of available qubits in the quantum processor. Using the outcomes measured during teleportation, we calculate and apply byproduct operators through dynamic circuits to correct local transformations induced on the teleported state. We evaluate the quality of teleportation by tracing the preserved entanglement and fidelity of the teleported two-qubit graph state from its density matrix. In the real-machine experiments, we demonstrate that 58% of the teleported state’s entanglement is sustained with a measured two-qubit negativity of $0.291\pm 0.018$ after three complete revolutions around the hamster wheel, or equivalently, after hopping across 56 qubits. On the machine-specific noisy emulator, we found that the teleported state after 100 hops still sustained 45% of its entanglement. By performing teleportation along a regenerating graph state, our work is a step forward in demonstrating the feasibility of measurement-based quantum computation.
@@
In the following contents, we cover the experimental setups and methodologies for implementing quantum hamster wheels, along with the mathematics of teleportation. Next, we introduce the concepts of negativity and fidelity as the two useful indicators of the amount of entanglement and the success of teleportation. Finally, we report and summarise the results obtained from the execution of quantum hamster wheels both on the emulator and on the real machine, which sustained the entanglement of the two-qubit state of high quality after 100 hops and 56 hops in teleportation, respectively.
@@
Figure 3: Polar plots of the teleported two-qubit graph state showing (a) negativity and (b) fidelity on the radial axis versus the number of teleportation hops in the angular axis sampled from the H1-1 emulator averaged over four trials. The results for larger than 75 hops are marked using different colours to highlight the glitch in the measured entanglement due to different noise parameters resulting from a long time gap between experiments. The three data points marked in red represent results obtained from the real H1-1 quantum computer for 9, 18 and 56 hops with respective negativities $0.459\pm 0.009,0.388\pm 0.014,0.291\pm 0.018$ and fidelities $0.958\pm 0.009,0.887\pm 0.014,0.791\pm 0.018$, where experiments were conducted on 21 March, 25 March and 24 April 2024 respectively. Assuming the distribution is Gaussian, the half-width of the shaded strip is two times the standard error of the sampled negativity or fidelity to represent 95% confidence interval. On the other hand, the data obtained from real H1-1 quantum computers are only sampled in a single trial due to limited resources. Therefore, we use the technique of bootstrapping to resample the data from the measured sample distribution for 200 trials. The error bars represent 95% confidence interval of the bootstrapped negativities or fidelities. All results are mitigated with REM as indicated in Teleportation quality witnesses. Due to the feedforward gates applied based on mid-circuit measurements in the dynamic circuit approach, REM is applied to only the two qubits of the final teleported state. The corresponding plot with regular crossing axes are included in Figure 4.
@@
In Figure 3a and Figure 3b, we present the negativities and fidelities recorded from the real machine (red-marked data points) and the emulator (blue data), plotted in polar plots against the number of hops that the two-qubit graph states were teleported across using the teleportation strategies. The emulated data is averaged over four trials. Due to limited access to computational resources, experiments on the H1-1 quantum computer were conducted for only three scenarios of 9, 18, and 56 hops in a single trial of the QST experiment with 1000 shots for each QST circuit. We bootstrap 200 samples from each QST basis measurement for both negativities and fidelities. The error bars of the data obtained from the H1-1 quantum computer of size $\epsilon$ take the half-range of the 95% confidence interval of the bootstrapped negativities or fidelities, such that $\epsilon=\frac{1}{2}{\left(\left\lfloor a_{0.975N}\right\rfloor-\left\lceil a_{0.025N}\right\rceil\right)}$, where $\{a_{n}\},n=1,2,\cdots,N$ is the sequence of bootstrapped negativities or fidelities ordered from lowest to highest and $N=200$ is the sample size. For both the negativity and fidelity, we have used REM, the nearest physical probability vector, and the nearest physical density matrix algorithm as previously outlined to mitigate classical readout errors. After all, from the H1-1 quantum computer we obtain the negativity of $0.459\pm 0.009,0.388\pm 0.014,0.291\pm 0.018$ and fidelity of $0.958\pm 0.009,0.887\pm 0.014,0.791\pm 0.018$ corresponds to 9, 18 and 56 hops respectively. For the data obtained from the emulator, we report a clear trend that the entanglement of the teleported state decreases gradually, almost linearly, as the number of hops increases. This trend is closely in agreement with the data points obtained from the real device without significant deviations, as well as the trend of the two-qubit gate fidelity plot in the H1 model product data sheet [16]. Therefore, it reinforces the reliability of the H1-1 emulator to approximate the corresponding device noise efficiently. There is a glitch that occurs at 76 hops, which results in a global increase in the measured negativity and fidelity afterwards. This is due to the long time gap between experiments performed before and after the 75-hop mark, where the noise parameters could have been optimised/totally different. From the plot, we also find that both dynamic circuit and post-selection approaches exhibit very similar behaviour regarding the amount of entanglement. This observation agrees with our expectations that the two approaches should only introduce little differences in the negativity and fidelity, given that their circuit depth differs by at most three gates. On the other hand, we have shown a distinct utility of the post-selection approach against dynamic circuits presented in previous work [6] using IBM quantum computers. This was a direct consequence of the accumulated latency for processing each feed-forward mid-circuit measurement sequentially and the relatively high level of noise in IBM quantum computers. From these results, we confirm that even after 56 hops, the negativity of the state teleported using a dynamic circuit can be preserved up to $0.291\pm 0.018$. When tested using the emulator, it still maintains a value of $0.224\pm 0.009$ up to 100 hops, which is just over 5 full cycles. Therefore, it is reasonable to expect that the actual limit of the number of hops the two-qubit state can teleport across using H1-1 quantum computer while preserving entanglement is far beyond 100.
@@
Discussion.—In summary, we demonstrated the teleportation of two-qubit graph states cyclically on a ‘quantum hamster wheel’. By making use of the reset gates, we were able to reuse the measured qubits and perform teleportation hops more than the total number of qubits available in the Quantinuum 20-qubit H1-1 quantum computer with ion-trap architecture. Despite limited real machine resources, in the emulation, we still showed that the entanglement of the two-qubit state was preserved with the negativity of $0.291\pm 0.018$ after 56 hops in teleportation on the real quantum processor and $0.224\pm 0.009$ after 100 hops on the corresponding emulator with the device-noise model. We expect this figure to be extended above 100 without the break-off of entanglement. Such performance benchmarks the utility of ion-trap quantum processors in achieving high-fidelity two-qubit gates and long coherence times that are more robust against perturbations. Moreover, since the teleportation protocol with dynamic circuits shares a similar principle and process as the single-qubit measurement-based quantum computation, these results also serve as a feasibility demonstration of measurement-based quantum computation.